OpenStudy (anonymous):

Solve: 6(x + 4) = 8(x − 2

OpenStudy (anonymous):

Solve: 6(x + 4) = 8(x − 2)

OpenStudy (anonymous):

like before first distrube the 6 and 8

OpenStudy (rc94):

Multiply brackets to get: 6x+24 = 8x -16, 40 = 2x, 20 = x

OpenStudy (anonymous):



OpenStudy (anonymous):

then u need to combine like terms

OpenStudy (anonymous):

oh i see thank you

OpenStudy (anonymous):

well get them on one side all x on one side and just numbers on the other

OpenStudy (anonymous):

its easier to place the x on the right and numbers on the left

OpenStudy (anonymous):

becuase of the negatives

OpenStudy (anonymous):

my teacher wants me to do this problem over the phone so thank you all for you help